Hi I have a question that I do not understand how to do. Would anybody please help me?

Use a proportion to estimate each population:


total deer counted: 64
tagged deer counted: 24
total deer tagged: 64

Can anybody explain how to answer this question please?Thanks

Answer by stanbon(75887)

This is a way of estimating the size of a wild population that
cannot conveniently be counted.
--------------
Here is the proportion usually used:
Total population / number caught
= tagged population released/tagged population caught later
------------
Your problem:
total deer counted: 64
tagged deer counted: 24
total deer tagged: 64
---------
Total population/ 64 = 64/24
Total pop = 64^2/24 = 170 2/3; rounded up is 171
==================
Cheers,
Stan H.